# Effective Field Neural Network

### v1

@@ Abstract
Figure 5: Continued function representation of an EFNN with 3 FP layers. (a) In an EFNN, the initial layer $S_{0}$ is recursively integrated into every quasi-particle layer. After a mapping, $S_{0}$ is multiplied with subsequent layers, forming a continued function representation (see right-hand side equation). (b) A typical architecture of ResNet, only one skip connection starts from $S_{0}$ and summation is used, which limits its renormalization capability. (c) Removing the $S_{0}$ connections transforms an EFNN into a standard DNN, where $S_{0}$ only appears at the beginning of the iterations, resulting in significantly diminished expressive power.
We give some discussions and conclude this paper. In both case studies, we have demonstrated that increasing the number of layers and neurons (or channels in convolutional layers) enhances the neural network’s accuracy, aligning with well-established findings in fields like computer vision. EFNN’s self-similar structure mirrors the renormalization process—a non-perturbative method for evaluating complicated functions, such as the Yang-Mills $\beta$ function [59] or Becke’s correction to the LDA functional in DFT [60]. This approximation framework stems from Padé approximants [61] using rational functions, which can also be expressed as continued fractions—an important and specific example of continued functions [62], known for superior convergence with much fewer parameters [63, 64, 65, 66]. Actually, EFNNs possess exactly the structure of continued functions, as shown in Fig. 5(a). In an EFNN, the initial feature layer $S_{0}$ connects recursively to every quasi-particle layer via mappings $g_{j}$; each nonlinear mapping $f_{j-1}$ is multiplied with $g_{j}(S_{0})$. In addition, EFNNs employ the $\tanh$ activation function within their renormalization-like structure, making them compatible with standard deep learning frameworks and enabling training via stochastic gradient descent with back-propagation. This contrasts with Padé approximants, which typically use the inverse function and hence usually determine parameters through perturbative series matching due to potential singularity in the gradient descent. As shown in Fig. 5(b), ResNets  [67, 68, 69, 70] adopt a similar but fundamentally different skip-like connections from EFNNs. In a ResNet, $S_{0}$ is skip-connected to later layers only once, with later skip connections starting in the middle, and its summation-based aggregation fails to meet the continued function formulation—resulting in suboptimal performance to characterize the many-body interactions. Standard DNNs (Fig. 5(c)) lack skip connections entirely, and $S_{0}$ only appears in the beginning of iterations, resulting in a very limited expressive capability. Moreover, by partitioning intermediate layers into effective field and quasi-particle layers, EFNNs provide a clear physical interpretation: quasi-particles emerge from the dot product of interacting particles (via a mapping) and effective fields, which is absent in ResNets and standard DNNs.
In conclusion, inspired by field theory, we have proposed a novel deep neural network architecture called EFNN, which effectively captures high-order interactions in both classical and quantum models. EFNN’s high accuracy and efficiency arise from the recursive refinement of effective fields and the construction of quasi-particles that encapsulate complex many-body effects. As discussed earlier, the connections from the true particle layer $S_{0}$ to the quasi-particle layers (or the preceding symmetrization layers) establish a structure of continued functions, and have renormalization power analogous to Padé approximants. Once the effective fields and quasi-particles are properly renormalized, the trained model can accurately predict the energy for a system with complicated many-body interactions. Furthermore, EFNNs maintain high accuracy even when applied to significantly larger systems. In future work, we aim to integrate Padé approximation and continued function theories more rigorously into the EFNN framework and derive error estimates for the neural network’s predictions.
@@
Acknowledgments—This work is supported by the Hong Kong Research Grants Council (16306220) and National Key R&\&&D Program of China (2021YFA1401500). Y.Z. is supported by the Max Planck Partner lab on quantum materials from Max Planck Institute Chemical Physics of Solids.

### v2

@@ Abstract
Figure 5: Comparison of neural network structures. (a) In an EFNN, the initial layer $S_{0}$ is recursively integrated into every quasi-particle layer. After a mapping, $S_{0}$ is multiplied with subsequent layers, forming a continued-function representation (see the equation above the figure). (b) A typical architecture of ResNet, only one skip connection starts from $S_{0}$ and summation is used, this is not a correct continued-function structure. (c) Removing the $S_{0}$ connections transforms an EFNN into a standard DNN, with no continued-function structure. (d) A typical architecture of DenseNet, $C$ is concatenation operation, this is not a correct continued-function structure.
We give some discussions and conclude this paper. In all three case studies—the classical 3-spin infinite-range model, continuous classical Heisenberg system, and quantum double exchange model—EFNNs consistently outperform DenseNets, ResNets, DNNs, and the effective model, achieving the lowest errors. Particularly striking is EFNN’s ability to generalize: trained on $10\times 10$ lattices, it accurately predicted energies for systems as large as $40\times 40$, while DenseNets, ResNets, and standard DNNs struggle with this extrapolation task. This superior performance isn’t accidental—it stems from EFNN’s mathematical foundation in continued functions. While conventional architectures like DenseNet and ResNet were designed through trial and error with various architectural innovations, EFNN’s structure directly implements continued functions, a mathematical framework that naturally captures many-body interactions up to infinite order. More importantly, for a perturbative expansion like Eq. (3), EFNN selects the most relevant terms via training, and expresses these terms in nonlinear continued function structure. Fig. 5 compares these structures, demonstrating that only EFNN (Fig. 5(a)) possesses the correct continued function structure. The Supplementary Material [51] provides a detailed mathematical explanation of why EFNN implements continued functions while ResNet and DenseNet do not. A special case of continued functions is the Padé approximant, which takes the form of a continued fraction and is ubiquitously used throughout physics and numerical methods for its rapid convergence, strong extrapolation capability, and compact representation [59]. Continued functions generalize continued fractions by replacing the inversion operation with generic nonlinear functions. Here, we use tanh activation functions instead of inverses, which makes EFNN trainable through standard backpropagation—something impossible with traditional Padé approximants due to gradient singularities. This simple substitution preserves the mathematical power of continued functions while enabling practical neural network training. Notably, EFNN actually implements self-similar approximation [60], a powerful mathematical framework that has been successfully applied across diverse fields including high energy physics [61, 62, 63, 64, 65], condensed matter physics [66, 67], quantum mechanics [68, 69, 64, 70, 71, 72, 60], quantum field theory [73, 74, 60, 75], and applied mathematics [76, 66, 77, 78, 79, 80, 81, 82, 83, 84, 85]. Previously, applying self-similar approximation required tedious manual calculations to obtain order by order matching; EFNN changes this by automating the process through neural network optimization. EFNN’s self-similar structure also naturally comes from renormalization group equation, as detailed in the Supplementary Material [51].
In conclusion, we have developed EFNN, a neural network architecture inspired by field theory that captures high-order interactions in both classical and quantum models. The key to EFNN’s success lies in how it recursively refines effective fields and constructs quasi-particles to encode complex many-body effects. This partitioning into field and quasi-particle layers provides clear physical interpretation while maintaining computational efficiency. As we showed, the connections from the true particle layer $S_{0}$ to the quasi-particle layers create a continued function structure with renormalization properties similar to Padé approximants. What EFNN actually implements is self-similar approximation—a powerful mathematical technique that has been widely used in physics but previously required manual calculations for each problem. Once properly trained, this architecture learns effective representations that accurately predict energies even for systems with complicated many-body interactions. Our tests on three diverse systems—the classical 3-spin infinite-range model, continuous classical Heisenberg spin system, and quantum double exchange model—demonstrate that EFNN consistently outperforms DenseNet, ResNet, and standard DNNs. Most remarkably, EFNNs trained on small systems improve accuracy on larger ones, opening possibilities for training on expensive DFT data from small lattices then inferring on much larger systems. Looking ahead, we plan to develop theoretical error bounds and apply EFNN to other challenging problems in physics.
DATA AVAILABILITY
The data that support the findings of this study are openly available in Zenodo at Ref. [86].
@@
Acknowledgments—This work is supported by the Hong Kong Research Grants Council (16306220) and National Key R&\&D Program of China (2021YFA1401500). Y.Z. is supported by the Max Planck Partner lab on quantum materials from Max Planck Institute Chemical Physics of Solids.